Solve the system using the substitution method. If the system has no solution or an infinite number of solutions, state this.
-0.7x -0.1y= -15
0.4x +0.1y= 0.6

Answer by Alan3354(69443)

-0.7x -0.1y = -15
0.4x +0.1y = 0.6 --> 0.1y = 0.6 - 0.4x
Sub for 0.1y in the 1st eqn
---------
-0.7x - (0.6-0.4x) = -15
-0.7x - 0.6 + 0.4x = -15
-0.3x - 0.6 = -15
3x + 6 = 150
x = 48
------
0.1y = 0.6 - 0.4x
y = 6 - 4x
y = -186